OpenStudy (anonymous):

Write the equation of the line that passes through (2,2) and (6,3) in standard form

OpenStudy (anonymous):

use \[(y-y1)=m(x-x1)\] m=\[(y2-y1)\div(x2-x1)\] use y1 and x1 must be the same coordinate, my suggestion is use (2,2) y2 is 4 and x2 is 6

OpenStudy (anonymous):

I need the actual answer

OpenStudy (anonymous):

y-2=\[y-2=(1\div4) \times (x-2)\] y=\[(1\div4)x -(5\div2)\]

OpenStudy (anonymous):

no its +3/2 instead of -5/2

OpenStudy (anonymous):

The choices are 4x+4y=6 x+4y=12 4x+4y=12 4x+4y=-12 x-4y=-6

OpenStudy (anonymous):

we cant give answers

OpenStudy (anonymous):

we r here to help not solve

OpenStudy (anonymous):

stuff2do gave u the answer basssically, all u have to do is plug in the numbers